Please help me solve this equation.
Use substitution, addition/subtraction, or linear combination to solve the linear system.
10r-9s=18
6s+2r=1

Answer by mananth(16949)

10r-9s=18.....................1
6s+2r=1.......................2
multiply equation 2 by -5 and it to equation 1
-30s-10r+10r-9s=-5+18
-39s=+13
s= 13/-39
s=-1/3
..
plug the value of r in equation 1
10r-9*-1/3= 18
10r+3=18
10r=15
r=15/10
r=3/2